# Advanced Sign Language Video Generation with Compressed and Quantized Multi-Condition Tokenization

### v1

@@ 1 Introduction
Sign language, as a visual language, serves as the primary communication medium for deaf individuals. Early research focused on Sign Language Recognition (SLR) [27, 39, 75], Translation (SLT) [3, 17, 36, 12], or Production (SLP) [49, 47, 51, 71, 68]. More recently, Sign Language Video Generation (SLVG) [48, 52, 41] has gained increasing attention, which aims to generate realistic and expressive sign language videos from spoken language texts, preserving the unique identity of a target signer, as shown in Figure 1(1). This growing interest is driven by its potential applications in accessibility technologies, educational tools, immersive communication systems, etc.
SLVG presents a challenging problem due to the lack of explicit spatial or temporal alignment between the input (i.e., the spoken language texts) and output (i.e., the sign language videos) modalities. To address this, current SVLG methods focus on leveraging synchronized auxiliary conditions as an intermediary to align these two modalities. As shown in Figure 1(2a), most of the existing SLVG methods [48, 52] leverage skeletal sequences as an intermediary to bridge a text-to-skeleton translation model (i.e., an SLP model) and a skeleton-to-video generation model. Because the generative model is only guided by a single coarse condition (e.g., skeleton), such single-conditional methods struggle with the naturalness and expressiveness of the generated videos, particularly in capturing facial expressions and figure movements, as illustrated in Figure 1(3).
Recent advancements in human video generation have shown that incorporating fine-grained conditions (e.g., dense pose [69] or 3D models [7, 67]) or leveraging multiple conditions (e.g., depth combined with optical flow [70]) can substantially improve generative fidelity. Inspired by these, we consider whether multiple fine-grained conditions can be introduced to enhance the quality and expressiveness of generated sign language videos. As shown in Figure 1(2b), one intuitive solution is to extend the single-conditional methods by considering multiple fine-grained conditions as intermediaries, where a multi-condition translation model can directly predict capable of achieving multiple fine-grained conditions. However, as shown in Figure 1(4), we observe that directly translating such attributes is challenging due to their high-dimensional nature and susceptibility to errors. This raises an important question: How can we overcome the challenges in the translation of multiple fine-grained conditions to further advance SLVG?
@@ 2 Related Works
Sign Language Video Generation. Sign Language Video Generation (SLVG) aims to generate identity-preserving sign language videos from spoken language texts. Early methods decompose the task into two consecutive sub-tasks [48, 52], which are the text-to-skeleton translation (i.e., SLP) and the skeleton-to-video generation. SignGAN [48] first employs a transformer [61] with a mixture density formulation to translate spoken language text to skeletal sequence. Then, a GAN-based [15] skeleton-conditioned human synthesis model is introduced to generate sign language videos. FS-Net [52] extends SignGAN by predicting the temporal alignment to a continuous signing sequence. Because the single condition focuses solely on capturing basic pose structures while neglecting fine-grained details, the generated videos tend to appear less natural and expressive. SignGen [41] seeks to address this limitation through a novel end-to-end pipeline that integrates multi-condition guidance, including optical flow, pose, and depth. However, we argue that SignGen suffers from the issue of training-inference inconsistency. Specifically, during training, SignGen leverages multiple conditions extracted from the target video to enhance the generative performance. Yet, since these ground-truth conditions are inaccessible during inference, SignGen fails to utilize them, resulting in subpar performance. In this paper, we aim to develop a framework that leverages multiple fine-grained conditions to enhance the quality and expressiveness of generated sign language videos.
@@
Human Video Generation. Human video generation has advanced significantly with the deep generative models. Early approaches, such as Pix2PixHD [65] and vid2vid [64], leveraged GANs [15] to generate realistic images and videos from the structured inputs. Several works have also explored the human pose generation, conditioning on the whole body [1, 33, 34], face [8, 28], or hands [58, 30]. However, GAN-based methods often suffer from mode collapse and optimization challenges. More recently, diffusion models [55, 23, 63] have emerged as a robust alternative, producing high-quality images and videos with greater stability. Most prior diffusion-based approaches rely on ControlNet [73] and OpenPose [5] to process each video frame independently, neglecting the temporal consistency and leading to the inevitable flickering artifacts. Pose-guided diffusion [53, 24, 62, 76, 54], a class of image-to-video models conditioned on pose, addresses this issue by generating temporally consistent human videos while preserving appearance fidelity. Furthermore, recent research shows that incorporating fine-grained conditions, such as dense pose [69] or 3D models [7, 67], or leveraging multiple complementary conditions, such as depth and optical flow [70], can significantly enhance generative fidelity. Building on these advancements, we aim to harness state-of-the-art diffusion-based methods alongside multiple fine-grained conditions to advance SLVG further.
@@ 3.1 Preliminary
Diffusion Models. As a class of the generative models, the diffusion models [55, 23] consists of two processes, which are the diffusion process and the denoising process, respectively. In the diffusion process, the Gaussian noise is iteratively added to degrade the input sample over $T$ steps until the sample becomes completely random noise. In the denoising process, a denoising model is used to iteratively generate a sample from the sampled Gaussian noise. When training, given an input sample $\mathbf{x}_{0}$ and condition $\mathbf{c}$, the denoising loss is defined as
@@
Among them, $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$ is the noisy sample at timestep $t\in[1,T]$, where $\alpha_{t}$ is a predefined scalar from the noise scheduler. $\boldsymbol{\epsilon}$ is the added noise. $\boldsymbol{\epsilon}_{\theta}$ is the denoiser with the learnable parameters $\theta$, which predicts the noise to be removed from the noisy sample. Latent Diffusion Models (LDMs) [46] stands out as one of the most popular diffusion models. It performs the two processes in the latent space, which is encoded by a Variational Auto-Encoder (VAE) [26, 45].
Finite Scalar Quantization. Finite Scalar Quantization (FSQ) [35] is a concise quantization technique to compress continuous values into discrete values. FSQ can be an alternative of Vector Quantization (VQ) [16] in VQ-VAE [60]. Compared with VQ, FSQ does not suffer from codebook collapse and does not need complex machinery to learn expressive discrete representations. Specifically, given a scalar $z\in\mathbb{R}$ from the encoded latent, the quantized discrete value by FSQ is
@@ 3.2 Overview
Due to the lack of explicit spatial or temporal alignment between $\mathcal{T}$ and $\mathbf{X}$, current SVLG methods focus on leveraging synchronized auxiliary conditions (e.g., skeletal sequence) as intermediary to align them. Such pipeline paradigm can be formulated as
@@
Figure 2: Framework of our SignViP for sign language video generation (SLVG).
@@ 3.4 Sign Video Diffusion Model
Sign Video Diffusion Model aims to generate sign language videos in a diffusion-based manner [55, 23] under the guidance of the reference image $\mathbf{x}_{\star}$ and the continuous multi-condition embeddings $\mathbf{e}_{0:F-1}$. Inspired by the previous works [24, 76], as shown in Figure 2(3), Sign Video Diffusion Model consists of three modules: Condition Guider, Denoising U-Net, and Reference Net. Condition Guider and Reference Net respectively encode the multi-condition embeddings $\mathbf{e}_{0:F-1}$ and the reference image $\mathbf{x}_{\star}$ to guide the Denoising U-Net.
Denoising U-Net is the backbone of the Sign Video Diffusion Model, which mirrors the architecture of Stable Diffusion (SD) v1.5 [46]. Each U-Net block includes a ResNet layer [19], a self-attention layer, and a temporal-attention layer [18]. The self-attention layer and the temporal-attention layer perform attention operation [61] along the spatial axes and the temporal axis, respectively.
@@
Reference Net [24] shares the same architecture of SDv1.5 and operates in parallel with the Denoising U-Net. The reference image $\mathbf{x}_{\star}$ is first encoded into the latent space by the VAE encoder $\mathcal{E}_{\text{VAE}}$, $\mathbf{z}_{\star}=\mathcal{E}_{\text{VAE}}(\mathbf{x}_{\star})$. The encoded reference latent $\mathbf{z}_{\star}$ is then fed into the Reference Net. The output feature of the self-attention layer in each block of the Reference Net is spatially concatenated with the input feature of the self-attention layer in the corresponding block of the Denoising U-Net.
@@
The inference starts from the sampled Gaussian noise. Then, the diffusion scheduler (e.g., DDIM [56]) is applied to generate images with multiple denoising steps. For each inference step, the noise prediction relies on Classifier-Free Guidance (CFG) [22]. Finally, the generated video is achieved from the latent by a VAE decoder $\mathcal{D}_{\text{VAE}}$.
@@ 3.6 Multi-Condition Token Translator
As shown in Figure 2(1), following previous works [71, 2], the spoken language text $\mathcal{T}$ is firstly encoded by CLIP text encoder [43] to obtain semantic embeddings, which serve as the initial input hidden states of the GPT-2 model [42]. Each output hidden state of the GPT model is decoded through multiple parallel prediction heads to generate all tokens of the corresponding frame simultaneously. On the input side, tokens belonging to the same frame are mixed to obtain unified input hidden states. Unlike methods that require dedicated modules for video length prediction [68] or rely on real video lengths [49], Multi-Condition Token Translator naturally determines the video generation endpoint by producing an “[EOS]" token.
During training, given the pre-trained multi-condition encoder and FSQ Autoencoder encoder, the produced tokens are considered the ground-truth. The cross-entropy loss is computed between the predicted tokens $\hat{\mathbf{b}}_{0:F}$ and the ground-truth tokens $\mathbf{b}_{0:F}$ [^1] Note that we use $F$ instead of $F-1$ here due to the inclusion of the additional endpoint token “[EOS]”.,
@@ 4.1 Experimental Settings
Datasets. We employ two sign language datasets for experiments. (1) RWTH-2014T [3] is a German sign language dataset. It comprises 8,257 sign language videos. The dataset is divided into 7,096 training samples, 519 validation samples, and 642 test samples. To align with the 8$\times$ downsampling rate of VAE, the frame size was resized from 260$\times$210 to 272$\times$224. (2) How2Sign [9] is an American sign language dataset. It includes 2,456 sign language videos. Using the provided timestamps, we segmented the videos to create a sentence-level dataset. This dataset consists of 31,128 training samples, 2,322 test samples, and 1,741 validation samples. The frame size is set to 512$\times$512.
@@
BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE
Ground-Truth 33.06 20.81 15.00 11.90 34.27 20.37 13.11 9.78 7.53 21.43
SignGAN [52] 17.41 7.93 4.67 3.16 19.64 10.66 4.62 2.92 1.97 11.76
w/ AnimateAnyone [24] 18.29 7.75 4.59 3.23 19.70 11.82 4.85 2.83 1.92 12.12
SignGen [41] 13.28 3.05 1.13 0.51 16.13 8.21 1.91 0.64 0.41 9.54
SignViP (Ours) 26.72 15.65 11.14 8.65 28.85 16.21 9.36 6.28 5.04 16.99
@@
BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE
ProTran [49] 17.96 8.99 5.64 4.07 20.97 14.57 7.47 4.59 3.42 17.32
Adversarial [47] 17.70 8.96 5.72 4.18 21.15 14.76 7.15 4.66 3.48 17.84
MDN [51] 18.06 9.30 6.06 4.52 21.44 14.94 7.54 5.10 3.67 18.21
MoMP [50] 20.55 10.98 7.02 5.14 23.75 16.57 8.47 5.38 4.16 19.38
SignGAN [52] 12.14 6.10 3.88 2.85 14.79 10.86 5.27 3.30 2.62 13.19
w/ AnimateAnyone [24] 12.36 6.23 4.01 2.97 14.93 10.97 5.36 3.39 2.67 13.36
SignGen [41] 10.42 2.42 0.89 0.38 12.68 8.67 1.79 2.41 1.36 8.69
SignViP (Ours) 21.94 10.06 6.32 4.61 22.67 17.35 8.28 5.41 4.42 18.23
Evaluation Metrics. To evaluate semantic consistency, we utilize the back-translation metrics following ProTran [49]. Specifically, we train an SLT model [4] to translate sign language videos or poses back into texts. The back-translated text is then compared with the ground-truth text with metrics of BLEU [38] and ROUGE-L [29]. To evaluate the video quality, we employ FID [21], CLIP-FID, FVD [59], and Identity Similarity (IDS). Among them, CLIP-FID is a variation of FID that utilizes CLIP [43] embedding as the frame’s embedding. IDS measures the identity consistency between generated and ground-truth videos. It calculates the cosine similarity of face embeddings extracted using YOLO5Face [40] and Arc2Face [37]. To further investigate the generative capability of video diffusion models in addition to FVD [59], we employ frame-level metrics including PSNR [11], SSIM [66], and LPIPS [74], leveraging their suitability in scenarios where ground-truth videos are temporally aligned with the generated videos. Additionally, we introduce Hand SSIM, which measures SSIM specifically in the hand region for a more precise evaluation of the hand quality.
Implementation Details. In Multi-Condition Token Translator, we utilize a multilingual version of the CLIP model [^2] https://huggingface.co/sentence-transformers/clip-ViT-B-32-multilingual-v1 to enable handling of multiple spoken language texts effectively. In FSQ Autoencoder, the encoder and decoder follow the architecture of their counterparts in VAE. Specifically, FSQ Autoencoder applies 4 latent channels, with each channel having a quantization level of 5. Together, this results in a total vocabulary size of 625, computed as $6255^{4}=$ 625 due to the combination of levels across all channels. In Sign Video Diffusion Model, both the Denoising U-Net and the Reference Net are initialized with Stable Diffusion v1.5333https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5. The temporal-attention layers in the Denoising U-Net are initialized from AnimateDiff [18]. The condition augmentation rate is set to 0.001. During inference, Sign Video Diffusion Model utilizes a guidance scale of 3.5 for CFG. Additionally, the number of inference steps is configured to 50.
Training Details. The training of the three stages are conducted on 4 NVIDIA RTX A6000 GPUs using Adam optimizer [25], with each stage consisting of 50,000 training steps. The batch sizes of stage I, II, and III are 2, 16, and 16. Their learning rates are 1e-5, 5e-5, and 1e-6.
@@
SignGAN [52] 547.90 167.70 1431.38 0.463 667.44 210.11 2766.97 0.5377
        w/ AnimateAnyone [24] 595.99 161.97 1330.54 0.462 679.41 215.05 2484.39 0.5330
SignGen [41] 644.06 184.66 1715.32 0.515 815.69 186.32 3538.49 0.539
@@
ControlNet [73] 556.63 0.784 19.50 0.137 0.483 427.22 0.826 21.32 0.116 0.657
AnimateAnyone [24] 365.42 0.794 20.06 0.121 0.505 293.18 0.821 21.54 0.103 0.663

### v2

@@ 1 Introduction
Sign language, as a visual language, serves as the primary communication medium for deaf individuals. Early research focused on Sign Language Recognition (SLR) [koller2017re, pu2019iterative, zhou2020spatial], Translation (SLT) [camgoz2018neural, guo2018hierarchical, orbay2020neural, gan2021skeleton], or Production (SLP) [saunders2020progressive, saunders2020adversarial, saunders2021continuous, yin2024t2s, xie2024g2p]. More recently, Sign Language Video Generation (SLVG) [saunders2020everybody, saunders2022signing, qi2024signgen, shi2024pose] has gained increasing attention, which aims to generate realistic and expressive sign language videos from spoken language texts, preserving the unique identity of a target signer, as shown in Figure 1(1). This growing interest is driven by its potential applications in accessibility technologies, educational tools, immersive communication systems, etc.
SLVG presents a challenging problem due to the lack of explicit spatial or temporal alignment between the input (i.e., the spoken language texts) and output (i.e., the sign language videos) modalities. To address this, current SVLG methods focus on leveraging synchronized auxiliary conditions as an intermediary to align these two modalities. As shown in Figure 1(2a), most of the existing SLVG methods [saunders2020everybody, saunders2022signing] leverage skeletal sequences as an intermediary to bridge a text-to-skeleton translation model (i.e., an SLP model) and a skeleton-to-video generation model. Because the generative model is only guided by a single coarse condition (e.g., skeleton), such single-conditional methods struggle with the naturalness and expressiveness of the generated videos, particularly in capturing facial expressions and figure movements, as illustrated in Figure 1(3).
Recent advancements in human video generation have shown that incorporating fine-grained conditions (e.g., dense pose [xu2024magicanimate] or 3D models [corona2024vlogger, wei2025magicportrait]) or leveraging multiple conditions (e.g., depth combined with optical flow [xue2024follow]) can substantially improve generative fidelity. Inspired by these, we consider whether multiple fine-grained conditions can be introduced to enhance the quality and expressiveness of generated sign language videos. As shown in Figure 1(2b), one intuitive solution is to extend the single-conditional methods by considering multiple fine-grained conditions as intermediaries, where a multi-condition translation model can directly predict capable of achieving multiple fine-grained conditions. However, as shown in Figure 1(4), we observe that directly translating such attributes is challenging due to their high-dimensional nature and susceptibility to errors. This raises an important question: How can we overcome the challenges in the translation of multiple fine-grained conditions to further advance SLVG?
@@ 2 Related Works
Sign Language Video Generation. Sign Language Video Generation (SLVG) aims to generate identity-preserving sign language videos from spoken language texts. Early methods decompose the task into two consecutive sub-tasks [saunders2020everybody, saunders2022signing], which are the text-to-skeleton translation (i.e., SLP) and the skeleton-to-video generation. SignGAN [saunders2020everybody] first employs a transformer [vaswani2017attention] with a mixture density formulation to translate spoken language text to skeletal sequence. Then, a GAN-based [goodfellow2014generative] skeleton-conditioned human synthesis model is introduced to generate sign language videos. FS-Net [saunders2022signing] extends SignGAN by predicting the temporal alignment to a continuous signing sequence. Because the single condition focuses solely on capturing basic pose structures while neglecting fine-grained details, the generated videos tend to appear less natural and expressive. SignGen [qi2024signgen] seeks to address this limitation through a novel end-to-end pipeline that integrates multi-condition guidance, including optical flow, pose, and depth. However, SignGen suffers from training-inference inconsistency, which leads to suboptimal results. In this paper, we aim to develop a framework that leverages multiple fine-grained conditions to enhance the quality and expressiveness of generated sign language videos.
@@
Human Video Generation. Human video generation has advanced significantly with the deep generative models. Early approaches, such as Pix2PixHD [wang2018high] and vid2vid [wang2018video], leveraged GANs [goodfellow2014generative] to generate realistic images and videos from the structured inputs. Several works have also explored the human pose generation, conditioning on the whole body [balakrishnan2018synthesizing, ma2017pose, men2020controllable, shen2023advancing], face [deng2020disentangled, kowalski2020config], or hands [tang2018gesturegan, liu2019gesture]. However, GAN-based methods often suffer from mode collapse and optimization challenges. More recently, diffusion models [sohl2015deep, ho2020denoising, shen2025boosting, wang2025ensembling, shen2025imagharmony, shen2025imaggarment] have emerged as a robust alternative, producing high-quality images or videos with greater stability. Most prior diffusion-based approaches rely on ControlNet [zhang2023adding] and OpenPose [cao2017realtime] to process each video frame independently, neglecting the temporal consistency and leading to the inevitable flickering artifacts. Pose-guided diffusion models [shen2024imagpose, hu2024animate, wang2024v, zhou2024realisdance, shen2025imagdressing, shen2025long] addresses this issue by generating temporally consistent human videos while preserving appearance fidelity. Furthermore, recent research shows that incorporating fine-grained conditions, such as dense pose [xu2024magicanimate] or 3D models [corona2024vlogger, wei2025magicportrait], or leveraging multiple complementary conditions, such as depth and optical flow [xue2024follow], can significantly enhance generative fidelity. Building on these advancements, we aim to harness state-of-the-art diffusion-based methods alongside multiple fine-grained conditions to advance SLVG further.
@@ 3.1 Preliminary
Diffusion Models. As a class of the generative models, the diffusion models [sohl2015deep, ho2020denoising] consists of two processes, which are the diffusion process and the denoising process, respectively. In the diffusion process, the Gaussian noise is iteratively added to degrade the input sample over $T$ steps until the sample becomes completely random noise. In the denoising process, a denoising model is used to iteratively generate a sample from the sampled Gaussian noise. When training, given an input sample $\mathbf{x}_{0}$ and condition $\mathbf{c}$, the denoising loss is defined as
@@
Among them, $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$ is the noisy sample at timestep $t\in[1,T]$, where $\alpha_{t}$ is a predefined scalar from the noise scheduler. $\boldsymbol{\epsilon}$ is the added noise. $\boldsymbol{\epsilon}_{\theta}$ is the denoiser with the learnable parameters $\theta$, which predicts the noise to be removed from the noisy sample. Latent Diffusion Models (LDMs) [rombach2022high] stands out as one of the most popular diffusion models. It performs the two processes in the latent space, which is encoded by a Variational Auto-Encoder (VAE) [kingma2013auto, rezende2014stochastic].
Finite Scalar Quantization. Finite Scalar Quantization (FSQ) [mentzer2023finite] is a concise quantization technique to compress continuous values into discrete values. FSQ can be an alternative of Vector Quantization (VQ) [gray1984vector] in VQ-VAE [van2017neural]. Compared with VQ, FSQ does not suffer from codebook collapse and does not need complex machinery to learn expressive discrete representations. Specifically, given a scalar $z\in\mathbb{R}$ from the encoded latent, the quantized discrete value by FSQ is
@@
Figure 2: Framework of our SignViP for sign language video generation (SLVG). (1) The spoken language text is translated into the multi-condition tokens by Multi-Condition Token Translator. (2) These tokens are decoded by FSQ Autoencoder into multi-condition embeddings, which are equivalent to the embeddings of multiple fine-grained conditions (i.e., fine-grained poses and 3D hands) generated by a multi-condition encoder. (3) The embeddings are injected into Sign Video Diffusion Model to guide the generation of sign language videos.
@@ 3.2 Overview
Due to the lack of explicit spatial or temporal alignment between $\mathcal{T}$ and $\mathbf{X}$, current SVLG methods focus on leveraging synchronized auxiliary conditions (e.g., skeletal sequence) as an intermediary to align them. Such pipeline paradigm can be formulated as
@@ 3.4 Sign Video Diffusion Model
Sign Video Diffusion Model aims to generate sign language videos in a diffusion-based manner [sohl2015deep, ho2020denoising] under the guidance of the reference image $\mathbf{x}_{\star}$ and the continuous multi-condition embeddings $\mathbf{e}_{0:F-1}$. Inspired by the previous works [hu2024animate, zhou2024realisdance], as shown in Figure 2(3), Sign Video Diffusion Model consists of three modules: Condition Guider, Denoising U-Net, and Reference Net. Condition Guider and Reference Net respectively encode the multi-condition embeddings $\mathbf{e}_{0:F-1}$ and the reference image $\mathbf{x}_{\star}$ to guide the Denoising U-Net.
Denoising U-Net is the backbone of the Sign Video Diffusion Model, which mirrors the architecture of Stable Diffusion (SD) v1.5 [rombach2022high]. Each U-Net block includes a ResNet layer [he2016deep], a self-attention layer, and a temporal-attention layer [guo2023animatediff]. The self-attention layer and the temporal-attention layer perform attention operation [vaswani2017attention] along the spatial axes and the temporal axis, respectively.
@@
Reference Net [hu2024animate] shares the same architecture of SDv1.5 and operates in parallel with the Denoising U-Net. The reference image $\mathbf{x}_{\star}$ is first encoded into the latent space by the VAE encoder $\mathcal{E}_{\text{VAE}}$, $\mathbf{z}_{\star}=\mathcal{E}_{\text{VAE}}(\mathbf{x}_{\star})$. The encoded reference latent $\mathbf{z}_{\star}$ is then fed into the Reference Net. The output feature of the self-attention layer in each block of the Reference Net is spatially concatenated with the input feature of the self-attention layer in the corresponding block of the Denoising U-Net.
@@
The inference starts from the sampled Gaussian noise. Then, the diffusion scheduler (e.g., DDIM [song2020denoising]) is applied to generate images with multiple denoising steps. For each inference step, the noise prediction relies on Classifier-Free Guidance (CFG) [ho2022classifier]. Finally, the generated video is achieved from the latent by a VAE decoder $\mathcal{D}_{\text{VAE}}$.
@@ 3.6 Multi-Condition Token Translator
As shown in Figure 2(1), following previous works [yin2024t2s, baltatzis2024neural], the spoken language text $\mathcal{T}$ is firstly encoded by CLIP text encoder [radford2021learning] to obtain semantic embeddings, which serve as the initial input hidden states of the GPT-2 model [radford2019language]. Each output hidden state of the GPT model is decoded through multiple parallel prediction heads to generate all tokens of the corresponding frame simultaneously. On the input side, tokens belonging to the same frame are mixed to obtain unified input hidden states. Unlike methods that require dedicated modules for video length prediction [xie2024g2p] or rely on real video lengths [saunders2020progressive], Multi-Condition Token Translator naturally determines the video generation endpoint by producing an “[EOS]" token.
During training, given the pre-trained multi-condition encoder and FSQ Autoencoder encoder, the produced tokens are considered the ground-truth. The cross-entropy loss is computed between the predicted tokens $\hat{\mathbf{b}}_{0:F}$ and the ground-truth tokens $\mathbf{b}_{0:F}$ [^1] Note that we use $F$ instead of F1F-1 here due to the inclusion of the additional endpoint token “[EOS]”.,
@@ 4.1 Experimental Settings
[…]
Evaluation Metrics. To evaluate semantic consistency, we utilize the back-translation metrics following ProTran [saunders2020progressive]. Specifically, we train an SLT model [camgoz2020sign] to translate sign language videos or poses back into texts. The back-translated text is then compared with the ground-truth text with metrics of BLEU [papineni2002bleu] and ROUGE-L [lin2004automatic]. To provide a more comprehensive evaluation of back-translated texts, we further employ the COMET [rei2022comet, rei2020comet] metric, which is specifically designed to predict human judgments of machine translation quality. COMET is widely used for machine translation tasks [guerreiro2023hallucinations, alves2024tower, he2024exploring] and is considered more suitable than BLEU and ROUGE. To evaluate the video quality, we employ FID [heusel2017gans], CLIP-FID, FVD [unterthiner2018towards], and Identity Similarity (IDS). Among them, CLIP-FID is a variation of FID that utilizes CLIP [radford2021learning] embedding as the frame’s embedding. IDS measures the identity consistency between generated and ground-truth videos. It calculates the cosine similarity of face embeddings extracted using YOLO5Face [qi2022yolo5face] and Arc2Face [papantoniou2024arc2face]. To further investigate the generative capability of video diffusion models in addition to FVD [unterthiner2018towards], we employ frame-level metrics including PSNR [fardo2016formal], SSIM [wang2004image], and LPIPS [zhang2018unreasonable], leveraging their suitability in scenarios where ground-truth videos are temporally aligned with the generated videos. Additionally, we introduce Hand SSIM, which measures SSIM specifically in the hand region for a more precise evaluation of the hand quality.
Implementation Details. In Multi-Condition Token Translator, we utilize a multilingual version of the CLIP model [^2] https://huggingface.co/sentence-transformers/clip-ViT-B-32-multilingual-v1 to enable handling of multiple spoken language texts effectively. In FSQ Autoencoder, the encoder and decoder follow the architecture of their counterparts in VAE. Specifically, FSQ Autoencoder applies 4 latent channels, with each channel having a quantization level of 5. Together, this results in a total vocabulary size of 625, computed as $5^{4}=625$ due to the combination of levels across all channels. In Sign Video Diffusion Model, both the Denoising U-Net and the Reference Net are initialized with Stable Diffusion v1.5333https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5. The temporal-attention layers in the Denoising U-Net are initialized from AnimateDiff [guo2023animatediff]. The condition augmentation rate is set to 0.001. During inference, Sign Video Diffusion Model utilizes a guidance scale of 3.5 for CFG. Additionally, the number of inference steps is configured to 50.
Training Details. The training of the three stages are conducted on 4 NVIDIA RTX A6000 GPUs using Adam optimizer [kingma2014adam], with each stage consisting of 50,000 training steps. The batch sizes of stage I, II, and III are 2, 16, and 16. Their learning rates are 1e-5, 5e-5, and 1e-6.
@@
BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE COMET BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE COMET
Ground-Truth 33.06 20.81 15.00 11.90 34.27 0.6157 20.37 13.11 9.78 7.53 21.43 0.5882
MoMP [momp] + ControlNet [zhang2023adding] 19.12 8.95 5.33 3.61 21.54 0.5033 12.53 5.59 3.48 2.31 13.72 0.5122
MoMP [momp] + AnimateAnyone [hu2024animate] 20.05 8.79 5.24 3.72 21.68 0.5091 13.65 5.82 3.39 2.25 14.15 0.5208
SignGAN [saunders2022signing] 17.41 7.93 4.67 3.16 19.64 0.4977 10.66 4.62 2.92 1.97 11.76 0.5104
w/ AnimateAnyone [hu2024animate] 18.29 7.75 4.59 3.23 19.70 0.4928 11.82 4.85 2.83 1.92 12.12 0.5135
SignGen [qi2024signgen] 13.28 3.05 1.13 0.51 16.13 0.4086 8.21 1.91 0.64 0.41 9.54 0.4127
SignViP (Ours) 26.72 15.65 11.14 8.65 28.85 0.5608 16.21 9.36 6.28 5.04 16.99 0.5524
@@
BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE COMET BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE COMET
Ground-Truth 30.99 18.36 12.83 9.87 31.02 0.5978 24.56 14.96 10.31 7.91 24.88 0.6250
ProTran [saunders2020progressive] 17.96 8.99 5.64 4.07 20.97 0.5091 14.57 7.47 4.59 3.42 17.32 0.5549
Adversarial [saunders2020adversarial] 17.70 8.96 5.72 4.18 21.15 0.5127 14.76 7.15 4.66 3.48 17.84 0.5618
MDN [saunders2021continuous] 18.06 9.30 6.06 4.52 21.44 0.5251 14.94 7.54 5.10 3.67 18.21 0.5685
MoMP [momp] 20.55 10.98 7.02 5.14 23.75 0.5466 16.57 8.47 5.38 4.16 19.38 0.5802
SignGAN [saunders2022signing] 12.14 6.10 3.88 2.85 14.79 0.5123 10.86 5.27 3.30 2.62 13.19 0.5673
w/ AnimateAnyone [hu2024animate] 12.36 6.23 4.01 2.97 14.93 0.5231 10.97 5.36 3.39 2.67 13.36 0.5315
SignGen [qi2024signgen] 10.42 2.42 0.89 0.38 12.68 0.4324 8.67 1.79 2.41 1.36 8.69 0.4413
SignViP (Ours) 21.94 10.06 6.32 4.61 22.67 0.5347 17.35 8.28 5.41 4.42 18.23 0.5738
@@
SignGAN [saunders2022signing] 547.90 167.70 1431.38 0.463 667.44 210.11 2766.97 0.538
        w/ AnimateAnyone [hu2024animate] 595.99 161.97 1330.54 0.462 679.41 215.05 2484.39 0.533
SignGen [qi2024signgen] 644.06 184.66 1715.32 0.515 815.69 186.32 3538.49 0.539
@@
ControlNet [zhang2023adding] 556.63 0.784 19.50 0.137 0.483 427.22 0.826 21.32 0.116 0.657
AnimateAnyone [hu2024animate] 365.42 0.794 20.06 0.121 0.505 293.18 0.821 21.54 0.103 0.663